# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v1

@@ 5.2 Galactocentric gradients of line widths
Considering the low temperature (Table 5) of the sources, the thermal broadening of the emission lines can be neglected ($\sim 0.1-0.2$ km s-1), therefore, the line width is a good tracer of the gas turbulence of the sources. The decreasing Galactocentric trends observed for the FWHM of the studied molecules suggest that the sources tend to have lower turbulence as the distance from the Galactic center increases. Such a result would be in line with an environment becoming more and more quiescent at increasing $R_{\text{GC}}$. The nearest source to the Sun, WB89-002, is the only outlier that deviates from these trends.
This result is also consistent with the study of Lin et al. (2025), where the velocity dispersion of low-$J$ 13CO lines has been studied towards molecular clouds in the low-metallicity part of the Galactic disk. A decreasing trend of the line width is observed with $R_{\text{GC}}$, and the sub-virial dynamics observed in the metal-poor clouds hint at how the gas turbulence is not sufficient to counter-balance the cloud self-gravity. As stated in Lin et al. (2025), in low-metallicity environments, the magnetic field may play a more dominant role, replacing gas turbulence in sustaining the clouds.
@@ 5.3 Galactocentric gradients of molecular ratios
Studies on the formation pathways of HCO+, H13CO+ (Klemperer 1970; Millar et al. 1991), as well as HCO, H2CO, and CH3OH (Watanabe & Kouchi 2002; Fuchs et al. 2009; Minissale et al. 2016; Rivilla et al. 2019; Jiménez-Serra et al. 2025), have shown that their formation is linked to the presence of CO. In contrast, c-C3H2 is formed from ”free” carbon (not bound in CO) (Thaddeus et al. 1985; Sipilä et al. 2016). Moreover, it is known that [C/O] decreases with the distance from the Galactic center (Méndez-Delgado et al. 2022). Therefore, we expect to have less ”free” carbon in the OG to produce c-C3H2, because most of it would be locked in CO. This would lead to a prediction of an increasing trend for the CO/c-C3H2 ratio.
Considering the lack of direct CO measurements for our sample, we can use species containing C and O, and formed from CO, such as HCO+, H13CO+, HCO, H2CO, and CH3OH. In Fig. 3, the Galactocentric gradients of these ratios are presented. No clear trends arise from the data, also characterized by $\left|\rho\right|<0.2$ (except for methanol, for which we have $\rho\sim-0.3$). The ratio between H13CO+ and c-C3H2 is the only one presenting a marginal increasing trend (in agreement with the formation pathways and the C/O abundance), however, too shallow to be considered reliable.
Figure 3: Galactocentric gradients of molecular ratios. The plot shows the trend of the ratios between CO molecule products (i.e., HCO+, H13CO+, HCO, CH3OH, and H2CO) and c-C3H2, as a function of Galactocentric radius ($R_{\text{GC}}$). The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The linear fit results are shown as the dark blue lines. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown.
@@ 5.4 Caveats and uncertainties
The comparison between observations and chemical models by Fontani et al. (2024) for WB89-670 suggests that the methanol could be sub-thermally excited: the kinetic temperatures predicted by chemical modeling are in the range $\sim 20-40\penalty 10000\$K, while the excitation temperatures are of the order of $\sim 10$ K. Therefore, an important caveat arises from the assumed $T_{\text{ex}}$ for the species for which we cannot estimate it directly from the observations.
Furthermore, the sources are located in a range of heliocentric distances, $d$, between 3 and 15.1 kpc, which implies that the linear scale probed is not the same in all targets. Fontani et al. (2024) demonstrates the presence of a high differentiation at smaller angular scales in WB89-670, and hence the average column densities derived in this work refer to linear scales with eventually different average conditions. However, this caveat is typical of source surveys, and in any case, we compare emissions arising from core envelopes in all targets.
Some results may also be influenced by low statistics, particularly within certain $R_{\text{GC}}$ ranges, where the data characterizing a few sources have a bigger impact on the gradients. This concern is more significant for formaldehyde and methanol, given their even lower sampling rates (only 18 molecular clumps, over 39, are observed in H2CO and CH3OH spectral windows), over a smaller $R_{\text{GC}}$ range (Fontani et al. 2022b).
Column densities of species such as HCN or HCO+ might be affected by large optical depth effects in the observed lines, which could impact the derived Galactocentric gradients. However, for HCN, the column densities were corrected by the line opacity (Table 7) obtained from the hyperfine structure of the observed transition, and both HCN and HCO+ were observed also in their less abundant 13C-bearing isotopologues (see Appendix E for the gradient of H13CN). Therefore, a gradient can also be inferred for these species using these isotopologues, expected to be optically thin.
Moreover, in the comparison between molecular and elemental trends, a further caveat comes from the fact that the elemental abundance gradients are extrapolated in the extreme OG ($R_{\text{GC}}\gtrsim 16$ kpc), where observations are sparse and thus the gradients are poorly constrained, and could be different from those at $R_{\text{GC}}\leq 16$ kpc. Indeed, recent studies have reported evidence for a flattening of the metallicity gradient at large Galactocentric radii, which could influence the observed molecular abundances and potentially explain the unexpectedly high chemical complexity in the outskirts of the Galaxy (Donor et al. 2020; Myers et al. 2022; Magrini et al. 2023; Da Silva et al. 2023).
@@ 6 Conclusions and outlooks
In this paper, we studied the chemical composition of 35 star-forming regions in the OG ($9\leq R_{\text{GC}}\leq 24$ kpc), using IRAM 30m telescope data from the CHEMOUT project. We identified 39 molecular clumps, and analyzed the molecular species detected in the greatest number of targets across our sample, i.e., HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). We estimated their column densities and calculated the fractional abundances of these species w.r.t. H2. We also updated the abundances of H2CO (18/18) and CH3OH (17/18) from Fontani et al. (2022b), using the H2 column densities obtained from the new NIKA2 continuum maps.
@@
The fractional abundances of the studied species with respect to H2 estimated assuming a gas-to-dust ratio dependent on $R_{\text{GC}}$ (Giannetti et al. 2017) show a decreasing trend with the Galactocentric radius. For HCN, HCO+, c-C3H2, HCO, SO, H2CO, and CH3OH the abundances scale as the elemental abundance of carbon ([C/H]) between the Solar circle and $\sim 24\penalty 10000\$kpc. This indicates that the efficiency in the formation of these molecules, which scales with the availability of the parent element (i.e., C), is at least as high as that found in the local Galaxy. Instead, the decreasing rate of SO is higher than that of [S/H], indicating a lower efficiency in the formation of this species in the OG. Conversely, the decreasing rate of H13CO+ presents the puzzling results of being lower than that of [13C/H], resulting in a higher formation efficiency in the OG, compared to the inner Galaxy.
@@
The Galactocentric gradients of the abundances indicate how the metallicity of the environment, which scales as well with $R_{\text{GC}}$, does not affect the formation efficiency of most of the molecules studied, even for so-called prebiotic molecules like formaldehyde. Metallicity appears to act mostly as a scaling factor for the molecular abundances relative to the elemental ones. The role of metallicity in the efficiency of molecule formation may become evident when studying molecules more complex than methanol, and/or it could be hidden by other environmental factors, such as the radiation field, gas temperature, or cosmic ionization rate, which may counterbalance its effect. These results confirm (after the previous papers of the CHEMOUT series) that the presence of organic molecules and tracers of protostellar activity is ubiquitous in the low metallicity environment of the OG.
@@
The next step for the CHEMOUT project is to infer a more precise kinetic temperature of the sources to better estimate the column densities from the line analysis. Determining this parameter will also help to reduce the space parameters on a chemical modeling analysis (e.g., using UCLCHEM; Holdship et al. 2017), which is crucial to infer the environmental differences of the star-forming regions in the different areas of the Galaxy, and understand which are the parameters driving the chemistry (Vermariën et al. 2025). The comparison between high-resolution observations and chemical models is crucial to disentangle the several contributions to these formation processes.
@@
It is furthermore crucial to enhance the statistics acquired as much as we can. In fact, given the scarcity of star-forming regions in the extreme OG, the data we have for $R_{\text{GC}}\gtrsim 20$ kpc have an important weight on the overall trends. Increasing the observed sample is essential to constrain the gradients obtained and to secure more precise results over the ongoing chemistry in this region of the Milky Way.
@@ Appendix A The CHEMOUT sample
Source aaaaaaThe sources labeled with a ”-b” are related to secondary velocity features detected in the spectra, likely due to a second molecular clump within the telescope beam. R.A. Dec $\ell$ $V_{\text{LSR}}$bbbbbbLocal standard of rest velocities estimated from the velocity peak of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, from Fontani et al. (2022). For the source WB89-315, being undetected in c-C3H2, we use the $V_{\text{LSR}}$ given in Blair et al. (2008). $R_{\text{GC}}$ccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of Russeil et al. (2017), and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in Hachisuka et al. (2015). Their kinematic Galactocentric distances are 15.7 and 18.9 kpc, respectively. ddccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of Russeil et al. (2017), and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in Hachisuka et al. (2015). Their kinematic Galactocentric distances are 15.7 and 18.9 kpc, respectively. $N_{\text{H}_{2}}$ddddddDerived from the NIKA2 continuum maps as described in Sect. 3.2. $T_{\text{ex}}$eeeeeeExcitation temperature of CH3OH, derived from the updated fits of the data presented in Fontani et al. (2022b) (see Sect. 4.1.1), adopted to fit the other species.
@@ Appendix B Millimeter continuum maps and H2 column densities
where $\nu_{2}$ and $\nu_{1}$, are 150 and 260 GHz (i.e., the frequencies corresponding to 2.0 and 1.2 mm), and $T_{\text{d}}$ is the assumed dust temperature (see Sect. 3.2). In Fig. 8, the Galactocentric trends of H2 are shown. The dataset estimated with a non-constant gas-to-dust ratio $\gamma=\gamma(R_{\text{GC}})$ (Giannetti et al. 2017) exhibits a positive trend with increasing values of $R_{\text{GC}}$, and a weak, positive correlation ($\rho=0.19$), which is due to the increasing $\gamma$ (Sect. 3.2).
@@
Figure 8: Gradients of column densities of H2, $N_{\text{H}_{2}}$, as a function of Galactocentric radius ($R_{\text{GC}}$). The light blue dataset represents the column densities estimated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue dataset shows the column densities calculated with a gas-to-dust ratio function of $R_{\text{GC}}$ from Giannetti et al. (2017). Both datasets have been fitted with a linear regression. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of the plot, the Pearson correlation coefficient, $\rho$, is shown (only for the dark blue dataset).
@@ Appendix C Line-fitting results
In this appendix, we present the results of the line fit from MADCUBA. Tables 7 and 8 present the best-fit parameters from MADCUBA for HCN, HCO+, c-C3H2, H13CO+, HCO, and SO. The fit results for H2CO and CH3OH have already been presented in Fontani et al. (2022b). The upper-limits values are also estimated for the undetected transitions. For the (self-) absorbed lines, no best-fit parameters are given.
In Figs. 9, 10, and 11, the observed spectra of the whole sample are shown with the Gaussian fits performed with MADCUBA for HCN, H13CO+, and SO. The plots showing the spectra of the remaining species are shown in Fontani et al. (2022) (c-C3H2, HCO+) and in Fontani et al. (2022b) (HCO, H2CO, CH3OH).
@@ Appendix E Galactocentric gradients of H13CN
Figure 13: Galactocentric gradients of fractional abundances w.r.t. H2, $X_{\text{mol}}$, of H13CN. The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from Giannetti et al. 2017). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The gradients of the elemental abundances of carbon (13C) and nitrogen (N), as reported by Méndez-Delgado et al. (2022), are plotted as dashed lines. The product of the parent elements of the species is represented by a black dash–dot line. All the elemental trends are plotted shifted to the start of the molecular gradients estimated with $\gamma(R_{\text{GC}})$ for reference to allow comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each plot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).

### v2

@@ 5.2 Galactocentric gradients of line widths
Considering the low temperature (Table 5) of the sources, the thermal broadening of the emission lines can be neglected ($\sim 0.1-0.2$ km s-1); therefore, the line width is a good tracer of the gas turbulence of the sources. The decreasing galactocentric trends observed for the FWHM of the studied molecules suggest that the sources tend to have lower turbulence as the distance from the Galactic center increases. Such a result would be in line with an environment becoming more and more quiescent at increasing $R_{\text{GC}}$. The nearest source to the Sun, WB89-002, is the only outlier that deviates from these trends.
This result is also consistent with the study of lin2025inadequate, where the velocity dispersion of low-$J$ 13CO lines has been studied toward molecular clouds in the low-metallicity part of the Galactic disk. A decreasing trend of the line width is observed with $R_{\text{GC}}$, and the sub-virial dynamics observed in the metal-poor clouds hint at how the gas turbulence is not sufficient to counterbalance the cloud’s self-gravity. As stated in lin2025inadequate, in low-metallicity environments, the magnetic field may play a more dominant role, replacing gas turbulence in sustaining the clouds.
@@ 5.3 Galactocentric gradients of molecular ratios
Studies on the formation pathways of HCO+, H13CO+ (klemperer1970carrier; millar1991gas), HCO, H2CO, and CH3OH (watanabe2002efficient; fuchs2009hydrogenation; minissale2016dust; rivilla2019first; jimenez2025modelling) have shown that their formation is linked to the presence of CO. In contrast, c-C3H2 is formed from ”free” carbon (not bound in CO) (thaddeus1985laboratory; sipila2016understanding). Moreover, it is known that [C/O] decreases with distance from the Galactic center (mendez2022gradients). Therefore, we expect to have less free carbon in the OG to produce c-C3H2, because most of it would be locked in CO. This would lead to a prediction of an increasing trend for the CO/c-C3H2 ratio.
Considering the lack of direct CO measurements for our sample, we can use species containing C and O that are formed from CO, such as HCO+, H13CO+, HCO, H2CO, and CH3OH. In Fig. 3, the galactocentric gradients of these ratios are presented. No clear trends arise from the data, which are also characterized by $\left|\rho\right|<0.2$ (except for methanol, for which we have $\rho\sim-0.3$). The ratio between H13CO+ and c-C3H2 is the only one presenting a marginally increasing trend (in agreement with the formation pathways and the C/O abundance); however, it is too shallow to be considered reliable.
Figure 3: Galactocentric gradients of molecular ratios. The plot shows the trend of the ratios between CO molecule products (i.e., HCO+, H13CO+, HCO, CH3OH, and H2CO) and c-C3H2, as a function of galactocentric radius ($R_{\text{GC}}$). The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The linear fit results are shown as the dark blue lines. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown.
@@ 5.4 Caveats and uncertainties
The comparison between observations and chemical models by fontani2024chemout for WB89-670 suggests that the methanol could be sub-thermally excited: the kinetic temperatures predicted by chemical modeling are in the $\sim 20-40\penalty 10000\$K range, while the excitation temperatures are on the order of $\sim 10$ K. Therefore, an important caveat arises from the assumed $T_{\text{ex}}$ for the species for which we cannot estimate it directly from the observations.
Furthermore, the sources are located in a range of heliocentric distances, $d$, between 3 and 15.1 kpc, which implies that the linear scale probed is not the same in all targets. fontani2024chemout demonstrated the presence of a high differentiation at smaller angular scales in WB89-670, and hence the average column densities derived in this work refer to linear scales with (eventually) different average conditions. However, this caveat is typical of source surveys, and in any case, we compare emissions arising from core envelopes in all targets.
Some results may also be influenced by low statistics, particularly within certain $R_{\text{GC}}$ ranges, where the data characterizing a few sources have a bigger impact on the gradients. This concern is more significant for formaldehyde and methanol, given their even lower sampling rates (only 18 molecular clumps of 39 are observed in H2CO and CH3OH spectral windows), over a smaller $R_{\text{GC}}$ range (fontani2022chemoutII).
Column densities of species such as HCN or HCO+ might be affected by large optical-depth effects in the observed lines, which could impact the derived Galactocentric gradients. However, for HCN, the column densities were corrected by the line opacity (Table 7) obtained from the hyperfine structure of the observed transition, and both HCN and HCO+ were observed also in their less abundant 13C-bearing isotopologues (see Appendix E for the gradient of H13CN). Therefore, a gradient can also be inferred for these species using these isotopologues, expected to be optically thin.
Moreover, in the comparison between molecular and elemental trends, a further caveat comes from the fact that the elemental abundance gradients are extrapolated in the extreme OG ($R_{\text{GC}}\gtrsim 16$ kpc), where observations are sparse and thus the gradients are poorly constrained, and could be different from those at $R_{\text{GC}}\leq 16$ kpc. Indeed, recent studies have reported evidence of a flattening of the metallicity gradient at large galactocentric radii, which could influence the observed molecular abundances and potentially explain the unexpectedly high chemical complexity in the outskirts of the Galaxy (donor2020open; myers2022open; magrini2023gaia; da2023oxygen).
@@ 6 Conclusions and outlooks
In this work, we studied the chemical composition of 35 star-forming regions in the OG ($9\leq R_{\text{GC}}\leq 24$ kpc), using IRAM 30 m telescope data from the CHEMOUT project. We identified 39 molecular clumps and analyzed the molecular species detected in the greatest number of targets across our sample; i.e., HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). We estimated their column densities and calculated the fractional abundances of these species with respect to H2. We also updated the abundances of H2CO (18/18) and CH3OH (17/18) from fontani2022chemoutII, using the H2 column densities obtained from the new NIKA2 continuum maps.
@@
The fractional abundances of the studied species with respect to H2 estimated assuming a gas-to-dust ratio dependent on $R_{\text{GC}}$ (giannetti2017galactocentric) show a decreasing trend with the galactocentric radius. For HCN, HCO+, c-C3H2, HCO, SO, H2CO, and CH3OH the abundances scale as the elemental abundance of carbon ([C/H]) between the solar circle and $\sim 24\penalty 10000\$kpc. This indicates that the efficiency in the formation of these molecules, which scales with the availability of the parent element (i.e., C), is at least as high as that found in the local Galaxy. Instead, the decreasing rate of SO is higher than that of [S/H], indicating a lower efficiency in the formation of this species in the OG. Conversely, the decreasing rate of H13CO+ presents the puzzling result that it is lower than that of [13C/H], resulting in a higher formation efficiency in the OG compared to the inner Galaxy.
@@
The galactocentric gradients of the abundances indicate how the metallicity of the environment, which scales as effectively with $R_{\text{GC}}$, does not affect the formation efficiency of most of the molecules studied, even for so-called prebiotic molecules such as formaldehyde. Metallicity appears to act mostly as a scaling factor for the molecular abundances relative to the elemental ones. The role of metallicity in the efficiency of molecule formation may become evident when studying molecules more complex than methanol, and/or it could be hidden by other environmental factors such as the radiation field, gas temperature, or cosmic ionization rate, which may counterbalance its effect. These results confirm (after the previous papers of the CHEMOUT series) that the presence of organic molecules and tracers of protostellar activity is ubiquitous in the low-metallicity environment of the OG.
@@
The next step for the CHEMOUT project is to infer a more precise kinetic temperature of the sources to better estimate the column densities from the line analysis. Determining this parameter will also help to reduce the space parameters on a chemical modeling analysis (e.g., using UCLCHEM; holdship2017uclchem), which is crucial to inferring the environmental differences of the star-forming regions in the different areas of the Galaxy and understanding the parameters driving the chemistry (vermarien2025understanding). The comparison between high-resolution observations and chemical models is crucial to disentangling the several contributions to these formation processes.
@@
It is furthermore crucial to enhance the statistics acquired as much as we can. In fact, given the scarcity of star-forming regions in the extreme OG, the data we have for $R_{\text{GC}}\gtrsim 20$ kpc weigh significantly on the overall trends. Increasing the observed sample is essential to constraining the gradients obtained and securing more precise results over the ongoing chemistry in this region of the Milky Way.
@@ Appendix A The CHEMOUT sample
Source aaaaaaThe sources labeled with a ”-b” are related to secondary velocity features detected in the spectra, likely due to a second molecular clump within the telescope beam. R.A. Dec $\ell$ $V_{\text{LSR}}$bbbbbbLocal standard of rest velocities estimated from the velocity peak of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, from fontani2022chemout. For the source WB89-315, being undetected in c-C3H2, we use the $V_{\text{LSR}}$ given in blair2008formaldehyde. $R_{\text{GC}}$ccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of russeil2017milky, and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in hachisuka2015parallaxes. Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. ddccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of russeil2017milky, and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in hachisuka2015parallaxes. Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. $N_{\text{H}_{2}}$ddddddDerived from the NIKA2 continuum maps as described in Sect. 3.2. $T_{\text{ex}}$eeeeeeExcitation temperature of CH3OH, derived from the updated fits of the data presented in fontani2022chemoutII (see Sect. 4.1.1), adopted to fit the other species.
@@ Appendix B Millimeter continuum maps and H2 column densities
where $\nu_{2}$ and $\nu_{1}$, are 150 and 260 GHz (i.e., the frequencies corresponding to 2.0 and 1.2 mm), and $T_{\text{d}}$ is the assumed dust temperature (see Sect. 3.2). In Fig. 8, the galactocentric trends of H2 are shown. The dataset estimated with a non-constant gas-to-dust ratio $\gamma=\gamma(R_{\text{GC}})$ (giannetti2017galactocentric) exhibits a positive trend with increasing values of $R_{\text{GC}}$, and a weak, positive correlation ($\rho=0.19$), which is due to the increasing $\gamma$ (Sect. 3.2).
@@
Figure 8: Gradients of column densities of H2, $N_{\text{H}_{2}}$, as a function of galactocentric radius ($R_{\text{GC}}$). The light blue dataset represents the column densities estimated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue dataset shows the column densities calculated with a gas-to-dust ratio function of $R_{\text{GC}}$ from giannetti2017galactocentric. Both datasets have been fitted with a linear regression. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of the plot, the Pearson correlation coefficient, $\rho$, is shown (only for the dark blue dataset).
@@ Appendix C Line-fitting results
In this appendix, we present the results of the line fit from MADCUBA. Tables 7 and 8 present the best-fit parameters from MADCUBA for HCN, HCO+, c-C3H2, H13CO+, HCO, and SO. The fit results for H2CO and CH3OH have already been presented in fontani2022chemoutII. The upper-limits values are also estimated for the undetected transitions. For the (self-) absorbed lines, no best-fit parameters are given.
In Figs. 9, 10, and 11, the observed spectra of the whole sample are shown with the Gaussian fits performed with MADCUBA for HCN, H13CO+, and SO. The plots showing the spectra of the remaining species are shown in fontani2022chemout (c-C3H2, HCO+) and in fontani2022chemoutII (HCO, H2CO, CH3OH).
@@ Appendix E Galactocentric gradients of H13CN
Figure 13: Galactocentric gradients of fractional abundances w.r.t. H2, $X_{\text{mol}}$, of H13CN. The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from giannetti2017galactocentric). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The gradients of the elemental abundances of carbon (13C) and nitrogen (N), as reported by mendez2022gradients, are plotted as dashed lines. The product of the parent elements of the species is represented by a black dash–dot line. All the elemental trends are plotted shifted to the start of the molecular gradients estimated with $\gamma(R_{\text{GC}})$ for reference to allow comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each plot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).